OpenStudy (anonymous):

How many roots does the polynomial function y=(x-8)(x+3)^2 have? A. 2 B. 3 C. 1 D. 4

OpenStudy (anonymous):

Let P(x) = 5x3 − 4x2 + 7x − 8. Then a root of that polynomial is 1 : P(1) = 5· 13 − 4· 12 + 7· 1 − 8 = 5 − 4 + 7 − 8 = 0. thats how u do it

OpenStudy (anonymous):

but not the answer

OpenStudy (anonymous):

Ok..

OpenStudy (welshfella):

this is a cubic equation ( x^3) and will have 3 roots

OpenStudy (welshfella):

This is the Fundamental Theorem of Algebra.

OpenStudy (anonymous):

its 2 so its A

OpenStudy (welshfella):

- that's not strictly true because 3 is a double root (x +3)^2

OpenStudy (anonymous):

im just telling you what it told me

OpenStudy (welshfella):

yes ok